OpenStudy (anonymous):

8 1/5* 7 3/4= estimate: exact:

OpenStudy (anonymous):

295.65

OpenStudy (anonymous):

estimate 8 1/5* 7 3/4= exact 8 1/5* 7 3/4=

OpenStudy (mertsj):

8 1/5 rounds to 8 7 3/4 rounds to 8 Estimate is 8 times 8 or 64

OpenStudy (mertsj):

\[\frac{41}{5}\times\frac{31}{4}=\frac{1271}{20}=63 \frac{11}{20}\]

OpenStudy (anonymous):

thanks you i got the 1271/20 too

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

how do you round?

OpenStudy (anonymous):

2 1/5 divided by 2 1/7

OpenStudy (mertsj):

If the fractional part is 1/2 or more round the whole number part to the next higher number. If the fractional part is less than 1/2 leave the whole number part the same.

OpenStudy (anonymous):

so 3 *3=9?

OpenStudy (mertsj):

Yes. 3 times 3 is 9

OpenStudy (anonymous):

sry so 10 1/3-2 7/9

OpenStudy (anonymous):

11*3?